# Structure from Noise: Confirmation Bias in Particle Picking in Structural Biology

### v2

@@ 6 Discussion and Outlook
Several tools in current cryo-EM pipelines aim to reduce bias by preprocessing the templates themselves [45]. For example, low-pass filtering the templates reduces their high-frequency features, which are most susceptible to overfitting by noise. This encourages the selection of broader structural motifs rather than fine-grained details, mitigating high-frequency confirmation bias. Our theoretical model (see Section 3) shows that this practice leads to reconstructions that converge toward low-pass-filtered versions of the templates. While not a complete remedy, such filtering is routinely employed in practice and is supported by empirical evidence, including the analysis presented in [5].
@@
In addition to template-based methods, several template-free approaches have been proposed that aim to detect candidate particles in an unsupervised manner, without relying on external references [23]. An alternative strategy within the template-based paradigm is to refine the similarity measures used during matching. Instead of standard cross-correlation, more robust alternatives, such as cross-entropy-based criteria [51], have been suggested. While promising, the development and validation of such metrics remain largely open challenges.
@@
It has been shown that below a certain SNR threshold, any particle-picking method is fundamentally doomed to fail [17]. A way to circumvent this limitation is to eliminate the particle-picking step altogether. Recently, it has been demonstrated, both theoretically and empirically, that one can reconstruct the 3D structure directly from raw micrographs at arbitrarily low SNR by bypassing particle picking and treating the unknown particle locations as nuisance variables [3]. Earlier direct-structure methods, such as those based on the method of moments [10] or on approximations to expectation-maximization [30], similarly aim to recover 3D volumes directly from micrographs without explicit particle selection. In principle, these approaches provide a clean way to avoid selection bias altogether.
@@ Acknowledgment
T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2019752, in part by ISF under Grant 1924/21, and in part by a grant from The Center for AI and Data Science at Tel Aviv University (TAD). We thank Amitay Eldar for insightful discussions and for his contributions to the initial formulation of the framework. T.B. thanks Amit Singer and Marc Gilles for bringing the phenomenon of hallucinated heterogeneity in conformational analysis to his attention.
@@
Appendix A provides the preliminaries for the appendix. Appendix B develops the theoretical framework and proofs for our general bias analysis of the GMM mean maximum-likelihood estimators, and its relation to the templates used in the template-matching selector. Appendix C extends this framework to the 3D reconstruction process in both single-particle cryo-EM and cryo-ET. In Appendix D, we describe in detail the empirical setup and procedures used for the simulations presented in Section 5. Appendix E compares the confirmation bias induced by template-matching selection, as analyzed in this work, with other sources of bias arising in single-particle cryo-EM and cryo-ET.
@@ C.1 Unified reconstruction model for cryo-EM and cryo-ET
Appendix D Setup of the numerical experiments
@@
Appendix E Different statistical models for confirmation bias in cryo-EM
Several stages of the cryo-EM and cryo-ET processing pipeline are vulnerable to confirmation bias, stemming from the use of initial templates or strong prior assumptions [52, Section 4.1]. Key examples include: (1) Particle picking, where 2D particles (in cryo-EM) or 3D subtomograms (in cryo-ET) are extracted using template-based or learning-based methods [46, 23, 12, 20]; (2) 2D classification (cryo-EM), where particles are clustered and averaged; (3) 3D reconstruction, including subtomogram averaging in cryo-ET, which is typically carried out using the expectation-maximization algorithm. As demonstrated and analyzed in [6], these algorithms are especially vulnerable to bias in low SNR conditions, where unverified assumptions can significantly shape the final reconstruction.
Each of these stages introduces bias through different statistical mechanisms. Figure 9 compares the forms of confirmation bias that occur during particle picking with those that emerge in downstream tasks, such as 2D classification. Below, we describe and contrast the Einstein from Noise phenomenon, which was presented in Section 1 and the confirmation bias observed in 2D classification with the bias arising from particle picking.
The Einstein from Noise phenomenon.
Although frequently mentioned in the context of template-based particle picking, the Einstein from Noise effect arises from a distinct model. In the Einstein from Noise experiment, all observations are aligned and averaged with respect to a fixed template. In contrast, particle picking involves selecting a small subset of patches from a large micrograph, specifically, those with the highest cross-correlation scores relative to the template. This key distinction gives rise to different types of confirmation bias and artifacts. Particle picking tends to produce reconstructions that closely resemble the template, differing only by a global scale factor (Theorems 3.1 and 4.2). In contrast, the Einstein from Noise effect produces an average that is positively correlated with the template but not identical [5].
In later stages of cryo-EM processing, such as 2D classification and 3D volume estimation, iterative algorithms often use predefined templates for initialization. When applied to data with no true signal (e.g., white Gaussian noise), these methods may still produce output that resembles the initial templates, even though the input data provides no evidence for such structure [4].
Figure 9(b) illustrates this phenomenon in the context of 2D classification. In this example, the input consisted entirely of pure noise, and the initial templates were set to $L=30$ distinct projections of the Beta-galactosidase structure, each of size $d=36\times 36$. A total of $M=5\times 10^{5}$ particles were processed using a 2D classification based on the expectation-maximization algorithm, run for 20 iterations. Even without a signal, the resulting class averages exhibit a positive correlation with the templates. However, unlike the case of template-based particle picking, the final class averages do not match the templates exactly (even up to a scale factor), though they still appear visually similar. It is worth noting that many modern 2D classification and 3D ab initio reconstruction algorithms now adopt reference-free initialization strategies (e.g., random initializations), which help mitigate such bias [27]. Nevertheless, when templates are introduced, whether explicitly or implicitly, similar confirmation effects can still occur.
Hallucinated heterogeneity in conformational analysis.
A distinct form of confirmation bias arises in methods for continuous heterogeneity analysis, particularly when state-of-the-art models are applied to structurally homogeneous datasets. In [22], the authors benchmarked RECOVAR [22], 3DFlex [41], and cryoDRGN [62] on a synthetic homogeneous dataset with high levels of noise. Despite the data consisting of a single conformation, all three methods showed signs of hallucinated heterogeneity. This demonstrates that both deep-learning methods (3DFlex and cryoDRGN) and classical statistical methods (RECOVAR) are not immune to this form of bias.

### v3

@@ 6.2 Strategies for mitigating confirmation bias
Several tools in current cryo-EM pipelines aim to reduce bias by preprocessing the templates themselves [43]. For example, low-pass filtering the templates reduces their high-frequency features, which are most susceptible to overfitting by noise. This encourages the selection of broader structural motifs rather than fine-grained details, mitigating high-frequency confirmation bias. Our theoretical model (see Section 3) shows that this practice leads to reconstructions that converge toward low-pass-filtered versions of the templates. While not a complete remedy, such filtering is routinely employed and is supported by empirical evidence, including the analysis presented in [5].
@@
A natural way to mitigate confirmation bias is to reduce or avoid structure-specific templates during particle picking. In practical cryo-EM workflows, this is often done by starting with template-free candidate generation, such as blob- or Laplacian-of-Gaussian (LoG)-based picking, followed by 2D classification. The resulting class averages can then be reused as internal references in later picking rounds. Unlike classical template matching with an external structural reference, blob/LoG methods rely only on weak, generic, low-resolution shape priors rather than molecule-specific templates. Accordingly, based on the results of this work, one expects these methods to favor coarse, blob-like, low-frequency structure rather than detailed structural features.
More broadly, several template-free approaches have been proposed for unsupervised particle detection without external references [21]. Such methods reduce the direct influence of externally imposed structural priors and therefore offer a potentially safer alternative to template-based picking. However, they are also expected to become unreliable below some SNR threshold.
A complementary mitigation strategy within the template-based paradigm is to refine the similarity measure used during matching. Rather than relying on standard cross-correlation, one may consider more robust criteria, such as cross-entropy-based measures [49]. While promising, the design, analysis, and empirical validation of such alternatives remain largely open challenges.
@@
Below a certain SNR threshold, any particle-picking method is fundamentally doomed to fail [17]. One way to circumvent this limitation is to eliminate particle picking altogether. Recently, it was shown, both theoretically and empirically, that one can reconstruct the 3D structure directly from raw micrographs at arbitrarily low SNR by bypassing particle picking and treating the unknown particle locations as nuisance variables [3]. Earlier direct-structure methods, including approaches based on the method of moments [10] and approximations to EM [28], likewise aim to recover 3D volumes directly from micrographs without explicit particle selection. In principle, these approaches avoid selection bias altogether.
6.3 Future work
Several directions remain open for future work.
Beyond the present picking model.
A natural next step is to extend the present theory beyond the pure-noise regime to the practically relevant low-SNR setting, where genuine particles are present. Such an analysis could quantify how genuine signal and template-induced bias interact in downstream reconstruction, and how it depends on the SNR, the density of the true particles, the picking threshold, and the mismatch between the true structure and the picking template.
Another important direction is to develop a theoretical framework for deep particle pickers. Unlike hand-crafted template-matching rules, deep pickers rely on learned scoring functions whose inductive biases are shaped by the training data, architecture, and optimization procedure. A rigorous analysis of such learned selection rules would broaden the scope of the present work and help clarify how modern data-driven picking pipelines may amplify or mitigate bias in low-SNR regimes.
A related extension is to move beyond controlled synthetic micrographs to realistic experimental cryo-EM data. Experimental micrographs contain structured background effects, such as ice gradients, carbon support edges, contamination, and variable ice thickness, that are absent from the idealized pure-noise model studied here.
CTF-aware modeling.
For analytical tractability, the present work does not model the CTF explicitly. In the pure-noise regime considered here, a fixed CTF acts as a linear filter, so Gaussian background noise remains Gaussian after CTF filtering, but with a modified covariance structure. From this viewpoint, CTF effects are partially absorbed into the correlated-noise model through the covariance matrix in Theorem 3.2. In practice, however, template matching is often CTF-aware and the CTF may vary across micrographs due to defocus changes. Extending the framework to explicit CTF-aware scoring and heterogeneous defocus is therefore an important direction for future work.
Cryo-ET validation and missing-wedge effects.
An important direction for future work is broader empirical and theoretical validation in a practical cryo-ET setting. While the present work includes controlled cryo-ET illustrations on synthetic pure-noise tomograms, these experiments are intended as proof-of-concept demonstrations of the bias mechanism rather than fully realistic subtomogram-averaging studies. Incorporating key experimental features such as missing-wedge effects [15], tomographic reconstruction artifacts, and validation on real cryo-ET datasets therefore remains for future work.
@@ Acknowledgment
T.B. is supported in part by BSF under Grant 2020159, in part by NSF-BSF under Grant 2019752, in part by ISF under Grant 1924/21, and in part by a grant from The Center for AI and Data Science at Tel Aviv University (TAD). We thank Amitay Eldar for insightful discussions and for his contributions to the initial formulation of the framework.
@@
Appendix A provides the preliminaries for the appendix. Appendix B develops the theoretical framework and proofs for our general bias analysis of the GMM mean maximum-likelihood estimators, and its relation to the templates used in the template-matching selector. Appendix C extends this framework to the 3D reconstruction process in both single-particle cryo-EM and cryo-ET. In Appendix D, we describe in detail the empirical setup and procedures used for the simulations presented in Section 5. Appendix 6.1 compares the confirmation bias induced by template-matching selection, as analyzed in this work, with other sources of bias arising in single-particle cryo-EM and cryo-ET.
@@ C.1 Unified reconstruction model for cryo-EM and cryo-ET
Appendix D Experimental methods
D.1 Setup of the numerical experiments
@@
D.2 Validation under stationary correlated Gaussian noise
To complement the white-noise experiments in the main text, we carried out an additional supplementary simulation study under stationary correlated Gaussian noise, with the goal of empirically validating the prediction of Theorem 3.2. This experiment should be viewed as the correlated-noise analogue of the white-noise validation presented in the main text: there, the selected pure-noise samples yield class centers that align with the templates themselves, whereas here the theory predicts that the selected class centers should align with a covariance-shaped transform of the templates.
Figure 11: Validation of template-induced bias under white and stationary correlated Gaussian noise. Top row: the original templates. Panel (a) shows the white-noise setting. In this case, the theoretical prediction for the normalized class centers coincides with the templates themselves, and the empirical class centers obtained from the selected noise samples closely match this prediction. Panel (b) shows the stationary correlated-noise setting. Here, the theory predicts that the normalized class centers are no longer equal to the original templates, but rather to their covariance-shaped transforms under the noise covariance operator, given by (D.3). The empirical class centers again closely agree with the theoretical prediction.
In this experiment, the input patches are drawn from a zero-mean stationary Gaussian field with a prescribed non-flat two-dimensional power spectrum. Specifically, if $W$ denotes white Gaussian noise and $S(k_{x},k_{y})$ is a nonnegative power spectrum, then each realization is generated as
$Y=\mathcal{F}^{-1}\!\big(\sqrt{S}\,\mathcal{F}(W)\big),$ (D.1)
where $\mathcal{F}$ and $\mathcal{F}^{-1}$ denote the two-dimensional discrete Fourier transform and its inverse, respectively. In the colored-noise experiment, we choose $S$ explicitly as a Gaussian low-pass spectrum,
$S(k_{x},k_{y})=\exp\!\left(-\frac{k_{x}^{2}+k_{y}^{2}}{2\ell^{2}}\right),$ (D.2)
where $\ell>0$ is a bandwidth parameter and $(k_{x},k_{y})$ range over the discrete Fourier frequencies. We then normalize $S$ by its average value so that the resulting spatial variance is approximately one. Equivalently, the covariance operator of $Y$ is a stationary convolution operator $\Sigma$, which acts by pointwise multiplication by $S$ in the Fourier domain. In contrast to the white-noise case, $\Sigma$ is not proportional to the identity, and therefore the induced template-matching bias depends explicitly on the correlation structure of the noise.
We then apply the same template-matching selection rule as in the main text to a large collection of pure-noise samples, using a bank of normalized templates $\{x_{\ell}\}_{\ell=1}^{L}$. For each selected sample, we record the template attaining the largest correlation score, and for each template class $\ell$ we compute the empirical conditional mean of the selected samples. Theorem 3.2 predicts that, after normalization by the selection threshold, this empirical class mean should align with
$\frac{\Sigma x_{\ell}}{x_{\ell}^{\top}\Sigma x_{\ell}}.$ (D.3)
In the white-noise case, where $\Sigma\propto I$, this expression reduces to the template itself. In the correlated-noise setting, however, it becomes a covariance-shaped transform of the template.
The results, shown in Figure 11, confirm this prediction. Panel (a) reproduces the white-noise benchmark: the theoretical and empirical normalized class centers both closely match the original templates. Panel (b) shows the stationary correlated-noise setting: although the theoretical outputs are now visibly distorted relative to the original templates, the empirical class centers remain in close agreement with the corresponding theoretical predictions. Quantitatively, this agreement is also reflected in a PCC close to one between the empirical and theoretical class centers. Thus, as in the white-noise case studied in the main text, template-based selection from pure noise still produces a structured biased output; the main difference is that under correlated noise the limiting bias is shaped not only by the templates, but also by the covariance operator $\Sigma$.